# for End-to-End Autonomous Driving

### v1

@@ 1 Introduction
Autonomous Driving (AD) technology has evolved rapidly in recent years, driven by advancements in artificial intelligence, sensor technology, and high-performance computing [1, 2, 3, 4]. However, real-world scenarios featuring unpredictable road users, dynamic traffic patterns, and diverse environmental conditions present significant challenges. Addressing these complexities requires sophisticated reasoning capabilities, allowing AD system to comprehend contextual information, anticipate user intentions, and make accurate real-time decisions. Traditionally, AD architectures have adopted a modular approach, with specialized components handling distinct aspects such as perception [5, 6, 7, 8, 9], mapping [6, 10], prediction [11, 12], and planning [13]. However, while this compartmentalization aids in debugging and optimizing individual modules, it often leads to scalability issues due to inter-module communication errors and rigid, predefined interfaces that struggle to adapt to new or unforeseen conditions [14, 15, 11, 16].
Recent advancements have seen the development of end-to-end systems that learn driving actions directly from sensor inputs, bypassing the need for symbolic interfaces and allowing for holistic optimization [17, 18, 19]. However, these systems, often being highly specialized and trained on narrow datasets, struggle to generalize effectively across diverse and complex real-world scenarios. This is where Multimodal Large Language Models (MLLMs) come into play, offering a transformative approach with their extensive training on wide-ranging datasets that encapsulate comprehensive world knowledge and advanced reasoning abilities through mechanisms like chain-of-thought reasoning [20]. Waymo’s proprietary EMMA model [1], built upon Google’s Gemini, exemplifies this trend, demonstrating significant advancements in integrating perception, decision-making, and navigation. Nevertheless, EMMA’s closed nature restricts access and experimentation for the wider research community.
@@
  • Finally, we fully release the codebase, datasets, and model weights utilized in OpenEMMA in https://github.com/dummy4submission/OpenEMMA for the research community to leverage, refine, and extend the framework, propelling further advancements in autonomous driving technology.
@@ 2.1 End-to-End Planning with Chain-of-Thought
Leveraging the power of pre-trained MLLMs [22, 23, 24, 25], we integrate the Chain-of-Thought reasoning process into the end-to-end trajectory planning process, following an instruction-based approach similar to that in [1]. Because MLLMs are trained with human interpretable knowledge, we prompt our MLLMs to also produce human interpretable knowledge. Therefore, unlike previous prediction methods that directly generate the trajectory in local coordinates [11, 3, 26], we instead generate two intermediate representations: the speed vector, $\mathbf{S}=\{s_{t}\}$, which denotes the magnitude of the vehicle’s velocity, and the curvature vector, $\mathbf{K}=\{k_{t}\}$, representing the turning rate of the vehicle. These presentations aim to reflect how a human drives: speed is how much the gas pedal should be pressed, whereas curvature is how much to turn the steering wheel.
@@ 3 Experiments
Method Model L2 (m) 1s L2 (m) 2s L2 (m) 3s L2 (m) avg Failure rate (%)
Zero-shot LLaVA-1.6 1.66 3.54 4.54 3.24 4.06
Llama-3.2 1.50 3.44 4.04 3.00 23.92
Qwen2-VL 1.22 2.94 3.21 2.46 24.00
OpeEMMA LLaVA-1.6 1.49 3.38 4.09 2.98 6.12
Llama-3.2 1.54 3.31 3.91 2.92 22.00
Qwen2-VL 1.45 3.21 3.76 2.81 16.11
Table 1: End-to-end trajectory planning experiments on nuScenes.
@@ 3.1 End-to-End Trajectory Planning
The experiments conducted on the validation set of the nuScenes dataset [21], and the models tested include GPT-4o [23], LLaVA-1.6-Mistral-7B [24] (refer to as LLaVA-1.6 for brevity thereafter), Llama-3.2-11B-Vision-Instruct [22] (refer to as Llama-3.2 for brevity thereafter), and Qwen2-VL-7B-Instruct [25] (refer to as Qwen2-VL for brevity thereafter). For comparison, we use the zero-shot method as the baseline, which relies solely on the historical ego status and the driving scene image, without incorporating any reasoning process. Furthermore, we set $T=5$, prompting the MLLM to predict the future trajectory over the next 5 seconds. Due to budget constraints and the need for reproducibility, the GPT-4o results are only conducted on a limited set of scenes and will be discussed in the case study.
Figure 2: YOLO 3D detection results. The class-color correspondences are: cars in gray, pedestrians in blue, traffic cones in orange, and trucks and trailers in brown. The blue-shaded surfaces indicate the heading of the object.
@@ 6 Limitation, and Future Work
Furthermore, due to the limited object grounding capabilities of current MLLMs, we incorporated a fine-tuned YOLO model into OpenEMMA to handle object detection tasks, rather than relying solely on the capabilities of the MMLM itself. While this approach provides a practical solution, it highlights the need for future advancements in MLLMs to bridge the gap in spatial reasoning and grounding accuracy. Addressing these limitations will be essential to achieve a truly unified framework that leverages MMLMs for all key perception and reasoning tasks in autonomous driving.

### v2

# OpenEMMA: Open-Source Multimodal Model for End-to-End Autonomous Driving

@@ 1 Introduction
Autonomous Driving (AD) technology has evolved rapidly in recent years, driven by advancements in artificial intelligence, sensor technology, and high-performance computing [18, 44, 59, 46, 13, 26, 23]. However, real-world scenarios featuring unpredictable road users, dynamic traffic patterns, and diverse environmental conditions present significant challenges [61]. Addressing these complexities requires sophisticated reasoning capabilities, allowing AD system to comprehend contextual information, anticipate user intentions, and make accurate real-time decisions. Traditionally, AD architectures have adopted a modular approach, with specialized components handling distinct aspects such as perception [65, 28, 21, 47, 17], mapping [28, 48], prediction [35, 45], and planning [51]. However, while this compartmentalization aids in debugging and optimizing individual modules, it often leads to scalability issues due to inter-module communication errors and rigid, predefined interfaces that struggle to adapt to new or unforeseen conditions [4, 20, 35, 43].
Recent advancements have seen the development of end-to-end systems that learn driving actions directly from sensor inputs, bypassing the need for symbolic interfaces and allowing for holistic optimization [16, 66, 29]. However, these systems, often being highly specialized and trained on narrow datasets, struggle to generalize effectively across diverse and complex real-world scenarios. This is where Multimodal Large Language Models (MLLMs) come into play, offering a transformative approach with their extensive training on wide-ranging datasets that encapsulate comprehensive world knowledge and advanced reasoning abilities through mechanisms like chain-of-thought reasoning [60, 37, 12]. Waymo’s proprietary EMMA model [18], built upon Google’s Gemini, exemplifies this trend, demonstrating significant advancements in integrating perception, decision-making, and navigation. Nevertheless, EMMA’s closed nature restricts access and experimentation for the wider research community.
@@
  • Finally, we fully release the codebase, datasets, and model weights utilized in OpenEMMA in https://github.com/taco-group/OpenEMMA for the research community to leverage, refine, and extend the framework, propelling further advancements in autonomous driving technology.
@@ 2.1 End-to-End Planning with Chain-of-Thought
Leveraging the power of pre-trained MLLMs [33, 36, 22, 57], we integrate the Chain-of-Thought reasoning process into the end-to-end trajectory planning process, following an instruction-based approach similar to that in [18]. Because MLLMs are trained to learn human-interpretable knowledge, we prompt them to produce outputs that remain interpretable. Consequently, unlike previous prediction methods that directly generate the trajectory in local coordinates [35, 59, 7], we instead generate two intermediate representations: the speed vector, $\mathbf{S}=\{s_{t}\}$, which denotes the magnitude of the vehicle’s velocity, and the curvature vector, $\mathbf{K}=\{k_{t}\}$, representing the turning rate of the vehicle. These presentations aim to reflect how a human drives: speed is how much the gas pedal should be pressed, whereas curvature is how much to turn the steering wheel.
@@ 2.2 Object Detection Enhanced by Visual Sepcialist
Method Model L2 (m) 1s L2 (m) 2s L2 (m) 3s L2 (m) avg Failure rate (%)
Zero-shot LLaVA-1.6-Mistral-7B 1.66 3.54 4.54 3.24 4.06
Llama-3.2-11B-Vision-Instruct 1.50 3.44 4.04 3.00 23.92
Qwen2-VL-7B-Instruct 1.22 2.94 3.21 2.46 24.00
OpenEMMA LLaVA-1.6-Mistral-7B 1.49 3.38 4.09 2.98 6.12
Llama-3.2-11B-Vision-Instruct 1.54 3.31 3.91 2.92 22.00
Qwen2-VL-7B-Instruct 1.45 3.21 3.76 2.81 16.11
Table 1: End-to-end trajectory planning experiments on nuScenes.
@@ 3.1 End-to-End Trajectory Planning
The experiments conducted on the validation set of the nuScenes dataset [6], and the models tested include GPT-4o [36], LLaVA-1.6-Mistral-7B [22], Llama-3.2-11B-Vision-Instruct [33], and Qwen2-VL-7B-Instruct [57]. For comparison, we use the zero-shot method as the baseline, which relies solely on the historical ego status and the driving scene image, without incorporating any reasoning process. Furthermore, we set $T=5$, prompting the MLLM to predict the future trajectory over the next 5 seconds. Due to budget constraints and the need for reproducibility, the GPT-4o results are only conducted on a limited set of scenes and will be discussed in the case study.
Figure 2: YOLO 2D detection results. The class-color correspondences are: cars in pink, trucks in green, trailers in yellow, pedestrians in blue, and traffic cones in white.
@@ 6 Limitation, and Future Work
Furthermore, due to the limited object grounding capabilities of current MLLMs, we incorporated a fine-tuned YOLO model into OpenEMMA to handle object detection tasks, rather than relying solely on the capabilities of the MLLM itself. While this approach provides a practical solution, it highlights the need for future advancements in MLLMs to bridge the gap in spatial reasoning and grounding accuracy. Addressing these limitations will be essential to achieve a truly unified framework that leverages MLLMs for all key perception and reasoning tasks in autonomous driving.